OpenStudy (anonymous):

A survey of 1,000 men and women asked, "Do you earn over $55,000 per year?" The table below shows the responses for males and females.

OpenStudy (anonymous):

Male, Female, Total Income over $55,000 475, 375, 850 Income below $55,000 75, 75, 150 Total 550, 450, 1,000 Based on these data, are "being male" and "earning over $55,000" independent events? (5 points) A. No, P(being male | the person earns over $55,000) ≠ P(being male) B. No, P(being male | the person earns over $55,000) = P(being male) C. Yes, P(being male | the person earns over $55,000) = P(being male) D. Yes, P(being male | the person earns over $55,000) ≠ P(being male)

OpenStudy (anonymous):

I think independence = true if P(A|B) = P(A) <--- or that A does not depend on B and if not true, then I guess we have not independent and I think P(A | B) = P(A) * P(B) if so then... let P(A) = P(person earns over $55k) P(B) = P(male) Then ask is this true, is P(person earns over $55k) * P(male) = P(male) ??? if so then we have independence. But Is P(male) = P(550 /1000) ??? or is P(male) in this case considered to be P(475/850) ? and is P(Person earns over 55k) = P(850/1000) ?

OpenStudy (kropot72):

Have you calculated the probability of a person being male, given that the person earns over $55,000?

OpenStudy (anonymous):

so P(475/850)

OpenStudy (anonymous):

and do we then compare that with just the probability of being male? P(550/1000) ?

OpenStudy (kropot72):

Correct. So the probability of a person being male, given that the person earns over $55,000 = 0.5588. The probability of a person being male is 0.550. When we look at the choices, the only correct one is A. Do you agree?

OpenStudy (anonymous):

looks good to me.. thank you. confusing, that my use of "P(male)" is ambiguous in a sense.. I must have the set notation confused I think.. will research it a bit and think about it. thank you

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

There are two tests for independence: \[P(A \cap B)=P(A) \times P(B)\] Also this test: \[P(A|B)=P(A)\] The second test is used for your question. As you can see the events in question are not independent

OpenStudy (anonymous):

So if I extract the numbers from the table P(A | B) is constructed as 475/850 The B part is taken from the total for the row, and the A part is taken from the male column. and means the probability of A given B, or probability of male given income of 55k. then the P(A) is constructed from the male column total / subject total, meaning probability of male given total number of people.

OpenStudy (kropot72):

Good work! You have understood :)